OpenStudy (anonymous):

Algebra 2 help please? I give medals. Let y be the difference of two numbers such that one number varies directly as x while the other number varies inversely as x. If y = –7 when x = –2, and y = 5 when x = 1, express y in terms of x.

OpenStudy (anonymous):

Let the numbers be a&b Then y=a-b,Let a>b also a=kx,where k is a constant. and\[b=\frac{ l }{ x }\] where l is a constant. when x=-2,a=-2k,\[b=\frac{ l }{ -2 },y=-7\] \[-7=-2k+\frac{ l }{ 2 },~or~-14=-4k+l ...(1)\] when x=1,a=k,b=l/1=l,y=5 ...(2) adding (1) & (2) -9=-3k k=3 5=3-l l=-2 a=3x \[b=\frac{ -2 }{ x }\] \[Hence~y=3x+\frac{ 2 }{ x }\]